# Neural models of multiscale systems: conceptual limitations, stochastic parametrizations, and a climate application

### v2

@@ Abstract
This work explores key conceptual limitations in data-driven modeling of multiscale dynamical systems, focusing on neural emulators and stochastic climate modeling. A skillful climate model should capture both stationary statistics and responses to external perturbations. While current autoregressive neural models often reproduce the former, they typically struggle with the latter. We begin by analyzing a low-dimensional dynamical system to expose, by analogy, fundamental limitations that persist in high-dimensional settings. Specifically, we construct neural stochastic models under two scenarios: one where the full state vector is observed, and another with only partial observations (i.e. a subset of variables). In the first case, the models accurately capture both equilibrium statistics and forced responses in ensemble mean and variance. In the more realistic case of partial observations, two key challenges emerge: (i) identifying the proper variables to model, and (ii) parameterizing the influence of unobserved degrees of freedom. These issues are not specific to neural networks but reflect fundamental limitations of data-driven modeling and the need to target the slow dynamics of the system. We argue that physically grounded strategies – such as coarse-graining and stochastic parameterizations – are critical, both conceptually and practically, for the skillful emulation of complex systems like the coupled climate system. Building on these insights, we turn to a more realistic application: a stochastic reduced neural model of the sea surface temperature field and the net radiative flux at the top of the atmosphere, assessing its stationary statistics, response to temperature forcing, and interpretability.
@@ I Introduction
A key challenge in current applied mathematics is to build data-driven models from partial observations of complex physical systems. Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1]. This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [2]. Seminal contributions include the work of Penland and collaborators [3, 4, 5], the multilevel modeling strategies of Kravtsov and colleagues [6, 7, 8, 9], and the contributions of Majda et al. [10, 11]. More recent proposal also include dynamical Markov-chain models to simulate the evolution of probability distributions as in [12, 13, 14] as well as novel ideas from generative models [15].
@@
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of the weather [16, 17, 18] and climate [19, 20, 21, 22, 23, 24, 25] systems. These tools build on earlier efforts in data-driven reduced-order modeling by considering new powerful architectures (i.e. neural networks) and by attempting to simulate the system in its full complexity and dimensionality. For example, [24] and [25] respectively seek to emulate the full ocean and atmosphere components of two state-of-the-art climate models. These models often excel at reproducing stationary statistics and generating forecasts across a range of time horizons, marking a promising direction for computationally efficient climate modeling. However, neural network emulators also exhibit several limitations. Long trajectory rollouts often suffer from error accumulation, which can drive the system toward instability. While promising approaches have been proposed to address this issue, e.g. [26, 27], these are currently research directions still to be explored across different models and scenarios. Instead current implementations maintain stability by prescribing part of the climate system as boundary conditions (e.g. [24] is an ocean emulator driven by pre-computed atmospheric fields as boundary conditions). We refer the reader to Pedersen et al. [26] for a detailed discussion of the stability challenges in autoregressive modeling. An additional issue, relevant for this study, is that skill in capturing unforced variability may not always guarantee accurate responses in perturbation experiments, limiting their use for causal mechanistic studies of climate dynamics. This leads to two important consequences: neural emulators (i) usually do not generalize in out-of-distribution climate change experiments and (ii) they may often not capture the right causal relations across variables. Note that the two issues (i) and (ii) above are likely related. The work of Van Loon et al. [28] provides us with a clear and insightful example of this limitation in the context of the “pattern effect” problem [29, 30], i.e. the influence of sea surface temperature (SST) warming patterns on the Earth’s energy imbalance. Van Loon et al. [28] considered the ACE2 emulator [21] and built a Green’s function by iteratively applying warming perturbations in different locations in the sea surface temperature (SST) field and computing the responses in the radiative flux at the top of the atmosphere (TOA). In doing so, [28] showed how such Green’s function does not capture in a quantitative way the causal relations between the SST and the radiative fluxes. This is just one example of the limitations of neural emulators in capturing responses to perturbations. Even if such issues are addressed in specific cases, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Other examples related to capturing forced responses to changes in the boundary conditions can be found in [24, 25]. Additionally, we note that current emulators usually focus on one climate component at a time, ocean or atmosphere, and the challenges described above, and also later in this paper, are expected to get worse when a fully coupled system will be considered. In general, limitations of neural emulators in capturing responses to external perturbations imply that such tools may not be yet ready to explore causal questions in climate dynamics.
@@
While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings. In the first part of this paper, we build on a series of previous works by A. Vulpiani and collaborators on the limitations of data-driven modeling [31, 32, 33]. We analyze a simplified stochastic dynamical system to illustrate fundamental limitations that are expected to persist in high-dimensional settings. Specifically, we ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations? In addressing this question, we consider two scenarios:
@@
  • Fully observed state vector. Data-driven modeling when the entire state vector is available. Here, we assume no knowledge of the underlying equations, but we have access to very long time series of the complete system state vector.
@@
  • Partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unknown. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
@@
To tackle such questions we proceed pedagocially by focusing on a “triad model” first proposed by Majda and collaborators in [34] and used as a testbed for Fluctuation-Dissipation theorem strategies in [35, 36]. The model mimics the abstract structure of dynamical cores in large-scale climate models with dynamics characterized by a linear operator and an energy conserving nonlinear, quadratic operator. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [37, 38]. Such operator can be retrieved from the model, and is here considered as the building block to get right for an emulator to, at least, capture the linear part of the response to general external perturbations. In the case of a fully observed state vector we show that the inferred neural model is able to reproduce stationary statistics of the system, i.e. the probability distributions and autocorrelation functions, as well as the response to small external perturbations. Therefore, there is no fundamental limitation for a neural model to respond to external perturbation, i.e. shifts in the input distribution, at least for perturbations of small amplitudes. However, in the case of partial observations, as in more realistic settings, we argue how the performance of a data-driven model necessarily depends on (i) the choice of variables to model and (ii) how to parametrize the unobserved degrees of freedom. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 39, 40, 41]. Therefore, we will argue that a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework, rather than as wholesale replacements.
@@
We emphasize that the focus on a simplified dynamical system is not a limitation in this context. In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings. As Smith (2002) [42] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models.
@@
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, where possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales. Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [40]. For an early instance of case (ii), see the work of Charney (1948) [43], Charney et al. (1950) [44] and the discussion in Dalmedico (2001) [41]. The example we consider simulates the joint evolution of the sea surface temperature (SST) field and the net radiative flux at the top of the atmosphere (TOA). Our motivation lies in understanding the spatial structure of climate feedbacks by addressing the so-called pattern effect, i.e. the dependence of Earth’s energy imbalance on the spatial patterns of surface warming [30]. To this end, we develop a simplified stochastic neural model based on 20 SST modes and a global mean, scalar time series for the radiative flux. Unobserved degrees of freedom are parametrized through multiplicative (state-dependent) noise. We demonstrate that the model captures both the stationary statistics and the radiative fluxes response to imposed temperature perturbations in two climate change experiments. Such model can then be used to explore the linkages between surface temperature warming patterns and radiative fluxes, as well as a “reconstruction” tool, by predicting the change in one variable given prescribed changes in the other. Finally, we apply ideas from response theory to infer direct causal relationships between surface temperature and TOA radiative flux. These results show qualitative agreement with perturbation experiments using atmospheric-only climate models and align with current physical understanding of radiative feedbacks in the climate system.
@@ II Theoretical Framework: Data-driven stochastic climate modeling and response theory
We begin by outlining the main theoretical concepts that underpin this work. First, we briefly introduce general ideas for constructing Langevin-like equations from partial observations of multiscale systems. We then review key aspects of response theory, which offers a general framework for understanding and assessing the ability of models to respond to perturbations.
@@
II.1 Reduced stochastic models: theoretical and practical considerations
Why reduced order modeling and effective equations?
Consider a nonlinear dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. The main interest in data-driven modeling, is to estimate the operator $\mathbf{f}$ directly from data. In doing so we face a serious limitation: in many high-dimensional complex systems we have access only to partial observations of the state vector $\mathbf{x}$. A temptation could be to use Taken’s embedding theorem [45]. While theoretically appealing, this strategy has significant practical limitations: it is not valid for stochastic systems and it becomes infeasible as the dimensionality of the system increases [46, 32, 40, 33, 47, 48]. A more promising direction leverages the multiscale nature of real-world dynamical systems by focusing on coarse-grained, effective dynamics. In some cases, it is possible to identify slow variables exhibiting low-dimensional effective dynamics and treat the cumulative effect of unresolved fast variables on the slow ones as stochastic forcing terms [49, 50, 51, 3, 4, 37, 36, 2, 52]. In these cases, the system state vector $\mathbf{x}$ can be decomposed into slow and fast components, $\mathbf{x}=(\mathbf{x}_{s},\mathbf{x}_{f})$, and motivates the derivation of an effective Langevin equation for the slow variables of the form:
$d\mathbf{x}_{s}=\mathbf{f}_{eff}(\mathbf{x}_{s})dt+\mathbf{\Sigma}(\mathbf{x}_{s})d\mathbf{W},$ (1)
where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}_{s}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process. The noise term parametrize the cumulative effect of unresolved fast processes $\mathbf{x}_{f}$ on the slow ones. In the general case the noise is “multiplicative”, i.e. state dependent: $\mathbf{\Sigma}(\mathbf{x}_{s})$. However, in many practical applications, additive noise can be preferred, with $\mathbf{\Sigma}$ independent of $\mathbf{x}_{s}$. In what follows, we will simplify notation by referring to $\mathbf{x}$ instead of $\mathbf{x}_{s}$, and we will clarify the precise meaning as needed in the text.
The reduced-order modeling approach emerges naturally in the study of large-scale climate dynamics, where recurrent spatiotemporal patterns, i.e. modes of variability, dominate the system’s behavior [53, 54]. Unlike traditional general circulation models (GCMs) that aim to simulate as many spatiotemporal scales as possible, reduced models target specific processes/problems, capturing a few key degrees of freedom while treating the rest as stochastic forcing. This not only reduces computational cost but also enhances mechanistic understanding, by isolating core processes and studying their qualitative behavior (see e.g. [55], as well as the introduction of [44]). Consequently, there is a growing need for systematic methods to reduce the dimensionality of high-dimensional systems and to quantify causal interactions across their degrees of freedom. This has fueled the development of low-dimensional stochastic models learned either from data [51, 3, 4, 56, 10, 7, 2] or from known equations [49, 50, 57, 58], as well as inference-based techniques to infer interactions across modes of variability [59, 60, 61, 62]. In this paper, we will primarily focus on data-driven stochastic modeling from partial observations. Nevertheless, we emphasize that even when all the relevant variables of the system are known, in multiscale physical systems there is often a need to formulate new effective (even deterministic) equations of the slow variables rather than focusing on the primitive equations. A classic example in the weather/climate literature is given by Charney (1948,1949) [63, 43] as well as Charney et al. (1950) [44]. Charney was among the first to recognize the importance of filtering out what he termed “meteorologically insignificant noise” from the primitive equations, leading to more interpretable and computationally efficient effective equations.
Data-driven reduced-order modeling: practical considerations and prior work.

### v3

# Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models

@@ Abstract
A central challenge in climate science and applied mathematics is developing data-driven models of multiscale systems that capture not only stationary statistics but also responses to external perturbations. Current neural climate emulators seek to resolve the atmosphere–ocean system in all its complexity but often fail to reproduce forced responses, limiting their use in causal studies such as Green’s function experiments. Response theory offers a rigorous framework to explore these limitations, moving beyond stationary statistics, probing causal mechanisms, and guiding model design. Using a simplified multiscale system, we argue that the skill of purely data-driven models in reproducing perturbed statistics depends critically on (i) identifying the proper variables to model, including spatial and temporal scales, and (ii) accurate stochastic parameterizations of unresolved processes. For studies of low-frequency climate dynamics, these insights motivate reduced-order neural models, tailored to specific processes, as valuable alternatives to general-purpose emulators. We illustrate this perspective with a reduced-order neural stochastic model designed to investigate responses of radiative fluxes to surface temperature perturbations. The model largely reproduces stationary statistics and enables the study of forced responses in both ensemble mean and variance, shifting the study of climate feedbacks from single trajectories to responses of probability distributions. These results underscore reduced-order models, augmented with modern neural networks and evaluated within the framework of response theory, as a practical strategy for causal inference and attribution studies in large-scale climate dynamics.
@@ I Introduction
A key challenge in current applied mathematics is to build data-driven models from partial observations of multiscale physical systems. Among the most challenging examples is the climate system: a forced, dissipative, and chaotic system whose dynamics span at least 15 orders of magnitude across spatial and temporal scales [1]. This problem has been extensively studied over the past three decades, mainly in the realm of reduced-order models [2]. Seminal contributions include the work of Penland and collaborators [3, 4, 5], the multilevel modeling strategies of Kravtsov and colleagues [6, 7, 8, 9], and the contributions of Majda et al. [10, 11]. Recent proposal include Markov-chain models as in [12, 13, 14] and novel ideas from generative models [15]. Most prior work has focused on evaluating reduced-order models on equilibrium statistics and forecasting, with limited attention to responses to external perturbations, an exception being the contributions of Majda and collaborators (see e.g. [16]). Capturing such responses is key for building models that not only mimic the system’s dynamics but also faithfully capture causal relations across degrees of freedom [17, 18].
@@
Recently, autoregressive models based on deep learning architectures have been proposed as potential emulators of complex turbulent systems such as the weather [19, 20, 21, 22, 23] and climate [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] systems. Unlike traditional reduced-order models, these neural architectures aim to emulate the full hierarchy of spatial and temporal scales present in the system. For example, [30] and [31] emulate the full ocean and atmosphere components of state-of-the-art climate models. Such emulators often reproduce stationary statistics and generate forecasts across a range of horizons, offering a promising direction for efficient climate modeling. However, several limitations remain. Long rollouts suffer from error accumulation that can drive instabilities [34, 35]. Moreover, current emulators are unable to capture extreme events if not already present in the training set, preventing their use for rare-event analysis in a changing climate [36]. More relevant here, skill in reproducing stationary statistics does not guarantee accurate responses in perturbation experiments, restricting their use for causal studies of climate dynamics. This has two consequences: (i) neural emulators often fail to generalize in out-of-distribution climate change scenarios, and (ii) they may not capture the correct causal relations across variables, two issues that are likely related. The work of Van Loon et al. [37] illustrates this limitation in the context of the “pattern effect” problem [38, 39], i.e. the influence of sea surface temperature (SST) warming patterns on Earth’s energy imbalance. Using the ACE2 emulator [26] trained on ERA5 [40], they constructed a Green’s function by perturbing SSTs regionally and measuring the radiative response. The resulting Green’s function failed to capture the causal relation between SST and radiative fluxes in a quantitative way. This is just one example of limitations of neural emulators in capturing responses to perturbations with additional examples in [30, 31, 41]. Even if such issues are addressed in specific cases, see [42] in the context of the pattern effect, they highlight a broader challenge: building general-purpose emulators that reliably capture forced responses remains inherently difficult and context-dependent. Moreover, most current emulators represent only a subsystem (e.g., atmosphere or ocean), and limitations in responding to perturbations are expected to worsen in coupled models. Overall, limitations in capturing forced responses suggest that neural emulators are not yet ready for causal applications in climate dynamics.
@@
While the field of neural emulation is advancing rapidly, it is useful to step back and examine its fundamental limitations in capturing forced responses, both from a theoretical standpoint and in controlled settings. In the first part of this paper, we build on a series of previous works by A. Vulpiani and collaborators on the limitations of data-driven modeling [43, 44, 45]. We consider a simplified stochastic dynamical system and ask the following question: Can neural autoregressive models, trained on (practically) infinitely long time series, accurately capture the response to general external perturbations? In addressing this question, we focus on two relevant scenarios in data-driven-only modeling:
@@
  • Unknown equations, fully observed state vector. Data-driven modeling when the entire state vector is available. Here, we assume no knowledge of the underlying equations, but we have access to very long time series of the complete system state vector.
@@
  • Unknown equations, partially observed state vector. Data-driven modeling in the presence of partial observations. This reflects common constraints in modeling high-dimensional systems from data, where the full set of relevant variables is typically unknown. In this more realistic setting, we again assume no knowledge of the underlying equations and only partial access to the system’s state vector.
@@
To tackle such questions we proceed pedagogically by focusing on a multiscale “triad model” first proposed by Majda and collaborators in [46, 47]. The model mimics the abstract structure of dynamical cores in large-scale climate models with a linear operator and an energy conserving nonlinear, quadratic operator. It provides a simplified yet relevant mathematical framework to unambiguously explore key limitations in capturing forced responses in climate emulators. We address the questions above by training neural stochastic emulators on a long integration of the triad model and test the models’ performance on stationary and perturbed cases. In doing so, we face the following issue: quantifying how an emulator respond to general forcing is an ill-posed question, as there are infinite possible forcings. Nonetheless, the linear response to a general forcing can be computed by convolving the forcing with an impulse response operator [48, 49]. This operator can be inferred from the model and is a building block that an emulator must capture to reproduce the response to small, general external perturbations [48], as well as the cause–effect relations among variables [17]. In the case of a fully observed state vector we show that the inferred neural model can reproduce stationary and perturbed statistics of the system. In the case of partial observations, as in realistic settings, the performance of a data-driven model necessarily depends on (i) the choice of proper variables to model and (ii) suitable parameterizations of unobserved processes. The field of reduced order modeling, introduced earlier, has long sought to confront these challenges by formulating effective Langevin equations capturing the coarse-grained, slow dynamics of the system [5, 50, 51], though typically without addressing responses to external perturbations. Therefore, we will argue that, in the case of low-frequency climate dynamics, a promising perspective is to build on such established approaches, emphasizing coarse-grained dynamics and stochastic parameterizations, while using neural networks as powerful tools within this framework.
@@
We emphasize that the focus on a simplified dynamical system is not a limitation in this context. In fact, the goal is not to demonstrate the success of a method, but to illustrate, by analogy, fundamental issues of data-driven modeling that are already evident in low dimensions and are likely to become more pronounced in higher-dimensional settings. As Smith (2002) [52] argues, while it is unreasonable to expect solutions to low-dimensional problems to generalize to high dimensions, so too it is unlikely that limitations observed in low-dimensional systems will vanish in high-dimensional, operational models. The first part of this paper is pedagogical in nature, aiming to bridge perspectives across researchers in statistical physics, climate science, and machine learning, while also providing practical methodological tools for the design and evaluation of data-driven models.
@@
Building on these ideas, we conclude this study by considering a real-world example. Throughout this paper, we emphasize, when possible, the distinction between “real-world examples/systems” and merely high-dimensional dynamics. By real-world examples we refer to situations where (i) we observe only the temporal evolution of a subset of variables without access to the full state vector (i.e. a projection of the full dynamics) or (ii) where even in the presence of all variables the system encompasses a huge number of spatial and temporal scales. Both cases (i) and (ii) lead to the need to formulate effective rather than primitive equations for both practical and conceptual reasons [50]. For an early instance of case (ii), see the work of Charney (1948) [53], Charney et al. (1950) [54] and the discussion in Dalmedico (2001) [51]. The example we consider is the “pattern effect” introduced above, which requires Green’s function experiments to quantify the response of top-of-atmosphere (TOA) radiative fluxes to perturbations in the sea surface temperature (SST) field. To this end, we develop a simplified stochastic neural model that emphasizes a large-scale representation of SST and global-mean net radiative flux, rather than attempting to resolve all spatial and temporal scales of variability as in current emulators. We discuss the coarse-graining and preprocessing choices relevant for model construction. The model largely reproduces both stationary and forced statistics and, importantly, enables Green’s function–like experiments to infer direct causal relations between SST and TOA radiative flux, a task where general-purpose emulators can perform poorly [37]. Crucially, we focus on responses in the ensemble mean and variance, rather than single trajectories as in most current studies, thereby extending the concept of climate feedback to include changes in distributional properties.
@@ II Theoretical Framework: Data-driven stochastic climate modeling and response theory
We begin by outlining the main theoretical concepts that underpin this work. First, we briefly introduce general ideas for reduced-order stochastic modeling relevant for climate dynamics. We then review key aspects of response theory, which offers a rigorous framework to move beyond the evaluation of stationary dynamics, probe causal mechanisms, and ultimately guide model design.
@@
II.1 Reduced-order stochastic climate models: theoretical and practical considerations
A general dynamical system can be represented in the form
$d\mathbf{u}/dt=\mathbf{F}+\mathbf{L}\mathbf{u}+\mathbf{n}(\mathbf{u}),$ (1)
where $\mathbf{F}$, $\mathbf{L}$ and $\mathbf{n}(\mathbf{u})$ represent an external forcing, a linear and nonlinear operators, respectively. $\mathbf{u}$ is not necessarily finite dimensional, and in the context of coupled atmosphere-ocean variability, it could in principle represent the deviation of all fields, at all scales, from their climatology [4]. However, for obvious practical reasons, Eq. (1) needs to be simulated on finite spatial and temporal resolutions. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is governed by a few dominant large-scale modes [4, 7, 48, 47, 55, 18]. The goal of reduced-order stochastic climate models is to develop low-dimensional models simulating large-scale, coarse-grained climate dynamics. The cumulative effect of unresolved fast variables, present at smaller spatial scales, is parametrized as stochastic terms. More formally, we decompose $\mathbf{u}$ into slow and fast components, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, leading to an effective Langevin equation of the form:
$d\mathbf{x}=\mathbf{f}_{eff}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W},$ (2)
where $\mathbf{f}_{eff}$ represents the deterministic “drift” dynamics, $\mathbf{x}$ are the slow variables and $\mathbf{W}$ is the standard Wiener process. The noise covariance matrix $\mathbf{\Sigma}$ can be independent of $\mathbf{x}$, corresponding to additive noise, or state-dependent $\mathbf{\Sigma}(\mathbf{x})$ corresponding to multiplicative noise. In the latter case, the equation is here interpreted in the Itô sense. Equation (2) is mathematically justified in the limit of infinite time-scale separation between fast and slow modes [1]. In the most general (and often intractable) setting, the effective dynamics of the slow variables also involve memory effects induced by the fast degrees of freedom, leading to non-Markovian dynamics. Kondrashov et al. (2015) [7] addressed this problem in a practical and efficient way by introducing the framework of multilevel stochastic models. Nonetheless, appropriate coarse-graining procedures have been shown to effectively average out non-Markovian contributions in many climate applications, yielding useful Markovian approximations [3, 4]. We refer the reader to [2] for a review on the role of non-Markovian contributions in effective climate equations and to [56, 57] for specific examples.
Reduced-order stochastic climate models, as in Eq. (2), provide an efficient alternative to state-of-the-art climate models for studying low-frequency climate variability. They are much more computationally efficient and allow us for mechanistic understanding of climate dynamics by isolating core processes and examining their qualitative behavior. This has fueled the development of low-dimensional stochastic models learned either from data [58, 3, 4, 16, 10, 7, 2, 59] or from known equations [60, 61, 62, 63, 46], as well as inference-based techniques to infer interactions across modes of variability [64, 65, 66, 18].